# Are Biological Systems More Intelligent Than Artificial Intelligence?

### v6

@@ 4 Formal Framework and Definitions
The core claim and Theorem 5.1 depend upon the Stack Theory primitives defined below. The surrounding narrative provides intuition. The ESM contains additional motivation, proofs, and related technical material [19, 15, 14, 16, 17, 12, 18].
Stack Theory describes an adaptive system in terms of the binary constraints or boundary conditions that define it. Embodiment is a choice of abstraction layer. This is a finite vocabulary of possible constraints (programs) representing the possible configurations of a body embedded in its environment [19]. These programs are used to express “statements” constraining the body. Policies and tasks are defined as statements in this embodied language, and delegation is captured by how task constraints are arranged across layers. ESM Note S5 provides a longer conceptual motivation and links to related perspectives.
For example, if a human raises their arm, then they are expressing a statement in their embodied language. That statement constrains what the body does next. The vocabulary $\mathfrak{v}$ of programs express this physical constraint. The inputs, outputs and policies of a system are all just subsets of its vocabulary. A $\mathfrak{v}$-task $\langle I,O\rangle$ is a set of inputs and corresponding outputs. If the $\mathfrak{v}$-task describes a ‘fit’ organism, then the outputs it contains are ‘correct’ according to natural selection. A $\mathfrak{v}$-task $\langle I,O\rangle$ can represent biological selection pressures by specifying which outputs are viable for which inputs. For example, an FEP-style agent can be treated as a task whose admissible futures minimise surprisal under an appropriate generative model [27, 39].
For readers who wish to check the formal dependencies (see also [19, 21]), the key definitions used in the theorem are summarised below.
@@
I set out to compare computers with biological systems. Since both are stacks of abstraction layers, I define multilayer architecture (MLA) below to describe them in comparable terms.
@@
  • The state of an MLA is a sequence of policies $\langle\pi^{0},\pi^{1},\dots,\pi^{n}\rangle$ and a sequence of vocabularies $\langle\mathfrak{v}^{0},\mathfrak{v}^{1},\dots,\mathfrak{v}^{n}\rangle$ such that for each $i<n$,
@@
I say the MLA is over-constrained when there exists $i<n$ such that $\Pi_{\lambda^{i}(\mathfrak{v}^{i})}=\emptyset$, and multilayer-causal-learning (MCL) is when the MLA is not over-constrained. Over-constraint means some layer cannot realise any policy compatible with the higher-level task. The MCL is the boundary where each layer still has at least one viable policy, so collective identity is possible.
@@ 5 Delegated Control and Collective Identity
The probability that the parts of a system embody a collective policy is higher if they embody weaker policies. This is because weaker policies are more likely to generalise between $\mathfrak{v}$-tasks [18].
Consider the aforementioned $\alpha_{1}$ and $\alpha_{2}$. A sufficiently weak policy at the level of individual cells is necessary for the collective to have an identity, because only the weakest policies in $\Pi_{\alpha_{1}}$ and $\Pi_{\alpha_{2}}$ will also be in $\Pi_{\omega}=\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$. A collective of cells remains cells regardless of whether I deem it to be an organ. The parts are still there. In terms of the stack, the cells $\alpha_{1}$ and $\alpha_{2}$ and the collective $\omega$ share a vocabulary. Just as the outputs of $\alpha_{1}$ can affect the inputs of $\alpha_{2}$ and vice versa, the outputs of $\omega$ determine the inputs of $\alpha_{1}$ and $\alpha_{2}$. The difference is a matter of scale. Just as w-maxing facilitates causal learning at the scale of an individual [14], in the context of an MCA it will facilitate learning at multiple scales, because the system as a whole is just another $\mathfrak{v}$-task.
@@ 5.1 The Stack
From the enactive perspective, a human is something cells do, rather than are. An organ boils down to a collection of cells. The cells are an abstraction layer in which the collective policy of the organ is expressed, like software. However, software is just a state of hardware [16]. Put another way, ‘organ’ is a label I ascribe to the behaviour of some cells, and ‘human’ is a label I ascribe to the behaviour of some organs. ‘x86’ is a label I ascribe to a product of human behaviour. It is a static abstraction layer an engineer has embodied in silicon and then left alone. ‘fire and forget’. Humans act as the lower layer from which x86 emerges, but once it is embodied in silicon we disconnect it from our stack. We causally ‘intervene’ in the stack. To understand the effect of this intervention, I need to abstract away the lower levels. I must “abstract away” cells and focus on the consequences of their collective behaviour.
@@
To construct layers I would start with a vocabulary $\mathfrak{v}^{i}$ at $i=0$, for example implied by a collective of cells or a lower level system. If this system exhibits goal directed behaviour, then it is given by the $\mathfrak{v}^{i}$-task $\lambda^{i}(\mathfrak{v}^{i})$ and a policy $\pi^{i}\in\Pi_{\lambda^{i}(\mathfrak{v}^{i})}$. $\pi^{i}$ implies behaviour. I translate policy into an abstraction layer using the aforementioned $\mathfrak{f}$. $\mathfrak{f}$ is applied to $\mathrm{Ext}(\pi^{i})$ to obtain a vocabulary $\mathfrak{v}^{i+1}$, and then $\lambda^{i+1}$ is applied to $\mathfrak{v}^{i+1}$ to obtain a task $\lambda^{i+1}(\mathfrak{v^{i+1}})$. If $\Pi_{\lambda^{i+1}(\mathfrak{v^{i+1}})}\neq\emptyset$, then there exists a common purpose or “self” at this level of abstraction and I select a another policy $\pi^{i+1}\in\Pi_{\lambda^{i+1}(\mathfrak{v^{i+1}})}$.
In narrative terms, adaptation at higher levels depends on adaptation at lower levels. I now state The Law of the Stack and give a brief proof sketch. The ESM provides a full proof and additional lemmas.
@@
Let $\langle\lambda_{0},\lambda_{1},\dots,\lambda_{n}\rangle$ be an MLA with abstractor $\mathfrak{f}$ and state $\langle(\pi_{0},\dots,\pi_{n}),(v_{0},\dots,v_{n})\rangle$ such that $v_{i+1}=\mathfrak{f}(v_{i},\pi_{i})$ and $\pi_{i}\in\Pi_{\lambda_{i}(v_{i})}$ for all $i<n$. Fix any $i<n$, and write
@@
Then the achievable utility at layer i+1i\!+\!1 is upper-bounded by the weakness of the policy at layer $i$ in the following sense.
@@
Utility $\epsilon(\gamma_{i+1})$ is attained by some correct policy at layer i+1i\!+\!1, whose extension is a subset of $L_{v_{i+1}}$, so $\epsilon(\gamma_{i+1})+|O_{\gamma_{i+1}}|\leq|L_{v_{i+1}}|$. Because $v_{i+1}$ is finite, $|L_{v_{i+1}}|\leq 2^{|v_{i+1}|}$. By construction of the abstractor, $v_{i+1}$ is the image of $\mathrm{Ext}(\pi_{i})$ under the map $o\mapsto T(o)$, hence $|v_{i+1}|\leq|\mathrm{Ext}(\pi_{i})|$. Combining yields $\epsilon(\gamma_{i+1})+|O_{\gamma_{i+1}}|\leq 2^{|\mathrm{Ext}(\pi_{i})|}$. ∎
@@ 5.2 Delegation and variational free energy
The Law of the Stack is a structural constraint. It says that degrees of freedom in higher layers is bounded by lower-layer slack. This aligns with the Free Energy Principle (FEP), which frames adaptive behaviour as variational free energy minimisation [39, 40, 71]. In the ESM (“Viability, lived history, and variational free energy”) I show that under a uniform viability prior, minimising free energy is equivalent to maximising the entropy (log-count) of viable future continuations. In Stack Theory terms, this is exactly w-maxing.
@@
Fix a planning horizon and let $\mu$ denote the viability task at layer i+1i{+}1. Write $O_{\mu}\subseteq L_{v_{i+1}}$ for the set of viable future continuations at that horizon. For any policy $\pi_{i+1}$ at layer i+1i{+}1, let $\Omega_{\pi_{i+1}}=\mathrm{Ext}(\pi_{i+1})\cap O_{\mu}$ be the set of viable futures compatible with $\pi_{i+1}$, and let $q_{\pi_{i+1}}$ be the uniform distribution on $\Omega_{\pi_{i+1}}$. Under the uniform viability prior assumed in the ESM, the corresponding (base-2) free energy satisfies
@@
In particular, free energy minimisation at layer i+1i{+}1 is bottlenecked by $|\mathrm{Ext}(\pi_{i})|$: reducing $F_{2}(q_{\pi_{i+1}})$ requires sufficient lower-layer slack, i.e. a weak enough realised policy at the lower layer.
@@ 5.3 Human Abstractions
If an abstraction layer is a window onto the world, then utility is how “good” the view is for a task. A static abstraction layer is a window that cannot move, while an MCA can change policy at lower levels, to move the window into an ideal location. When I create an ISA I insert a static abstraction layer at an intermediate level, so the vocabulary at the highest levels is limited to what I can do with those abstractions I have embodied in the ISA. In comparison, an MCA can learn policies at lower levels, to change the vocabulary of higher levels and facilitate multilayer causal learning. This shows how adaptability at higher levels demands adaptability at lower levels. I am not suggesting lower levels only optimise for utility at higher levels. After all, an increase in utility may increase the computational cost of searching for correct policies. Furthermore, a spatially extended environment constrains the size of vocabularies at all levels [17], because bounded systems can contain only finite information [8]. I mean lower levels learn an inductive bias for higher levels, balancing utility with other factors in service of goals at every scale.
@@ 6 Cancer-Like Failure Modes
I have explored some advantages of bio-mimetic distribution and delegation. Are there disadvantages? Yes. A state analogous to cancer exists for distributed systems with insufficient delegation in adverse conditions. Some have argued cancer is a loss of collective identity [57, 30]. In a biological system, when cells are isolated from the informational structure of their collective, they lose their high level identity. When that happens, they independently pursue their own lower level goals. They revert to primitive transcriptional behaviour, while the collective of which they were formerly part continues on without them. They escape from immune control mechanisms by becoming less differentiated, and closer to embryonic states.
@@
Proof. If some $\pi$ were feasible for all parts then $\pi\in\bigcap_{j=1}^{m}\widehat{\Pi}_{\alpha_{j}}=\widehat{\Pi}_{\omega}$, contradicting $\widehat{\Pi}_{\omega}=\emptyset$. Hence every realised policy fails at least one part’s constraints, so $S(\pi)\neq\{1,\dots,m\}$ for all $\pi\in L_{v}$. If $\widehat{\Pi}_{\alpha_{j}}\neq\emptyset$ then $\{j\}$ is a splinter since $\bigcap_{k\in\{j\}}\widehat{\Pi}_{\alpha_{k}}=\widehat{\Pi}_{\alpha_{j}}\neq\emptyset$. $\square$
@@
When $\Pi_{\lambda^{i}(v^{i})}=\emptyset$, the collective has no viable way to act as a coherent unit. Any continued function therefore requires dropping constraints, parent task splintering into ever narrower child tasks until those tasks are narrow enough to admit a policy. Part of the collective stops participating in the collective policy and reverts to a lower-level policy set with a much larger extension (more autonomy). This is a form of “selective forgetting” [14] applied to constraints rather than data.
This is analogous to cancer. $\alpha$ is isolated and can now pursue goals which do not align with those of the collective. This suggests systems should be minimally constrained so as to ensure graceful degradation or ‘sloppy fitness’ [2], maintaining core functionality in as wide a variety of unanticipated circumstances as possible. The more difficult the environment or the more top-down the system is controlled, the more tightly constrained behaviour becomes until the system becomes over-constrained and splinters. Co-operation becomes impossible and the parts of a system start to compete. Conversely, if control of a system is delegated to the lowest levels, then adverse conditions are less likely to eliminate all correct policies. Put another way, there will be less incentive for the parts of a system to splinter, as that would not greatly increase the number of correct policies available to them. A state analogous to cancer occurs in a distributed system when there is an over-reliance on top-down causation to make the parts of a system conform to a collective identity they are ill suited to, under the circumstances.

### v7

@@ 4 Formal Framework and Definitions
The core claim and Theorem 5.1 depend upon the Stack Theory primitives defined below. The surrounding narrative provides intuition. The ESM contains additional motivation, proofs, and related technical material [bennett2025thesis, bennett2023b, bennett2023c, bennett2024a, bennett2024b, bennettwelshciaunica2024, bennett2025e].
Stack Theory describes an adaptive system in terms of the binary constraints or boundary conditions that define it. Embodiment is a choice of abstraction layer. This is a finite vocabulary of possible constraints (programs) representing the possible configurations of a body embedded in its environment [bennett2025thesis]. These programs are used to express “statements” constraining the body. Policies and tasks are defined as statements in this embodied language, and delegation is captured by how task constraints are arranged across layers. ESM Note S5 provides a longer conceptual motivation and links to related perspectives.
For example, if a human raises their arm, then they are expressing a statement in their embodied language. That statement constrains what the body does next. The vocabulary $\mathfrak{v}$ of programs express this physical constraint. The inputs, outputs and policies of a system are all just subsets of its vocabulary. A $\mathfrak{v}$-task $\langle I,O\rangle$ is a set of inputs and corresponding outputs. If the $\mathfrak{v}$-task describes a ‘fit’ organism, then the outputs it contains are ‘correct’ according to natural selection. A $\mathfrak{v}$-task $\langle I,O\rangle$ can represent biological selection pressures by specifying which outputs are viable for which inputs. For example, an FEP-style agent can be treated as a task whose admissible futures minimise surprisal under an appropriate generative model [ciaunica2023b, Friston2010].
For readers who wish to check the formal dependencies (see also [bennett2025thesis, bennett2024appendices]), the key definitions used in the theorem are summarised below.
@@
I set out to compare computers with biological systems. Both are stacks of abstraction layers. Hence I define a stack (meaning a multilayer architecture or MLA) below. I do this to describe biological and contemporary computer systems in comparable terms.
@@
  • The state of the stack (MLA) is a sequence of policies $\langle\pi^{0},\pi^{1},\dots,\pi^{n}\rangle$ and a sequence of vocabularies $\langle\mathfrak{v}^{0},\mathfrak{v}^{1},\dots,\mathfrak{v}^{n}\rangle$ such that for each $i<n$,
@@
I say the stack (MLA) is over-constrained when there exists $i<n$ such that $\Pi_{\lambda^{i}(\mathfrak{v}^{i})}=\emptyset$, and multilayer-causal-learning (MCL) is when the stack is not over-constrained. Over-constraint means some layer cannot realise any policy compatible with the higher-level task. The MCL is the boundary where each layer still has at least one viable policy, so collective identity is possible.
@@ 5 Delegated Control and Collective Identity
The probability that the parts of a system embody a collective policy is higher if they embody weaker policies.
Classical cybernetics characterises regulation in terms of variety. Ashby’s law of requisite variety says that a regulator must possess at least as much variety as the disturbances it must absorb. Delegated control can be read as a constructive way to generate such requisite variety by distributing adaptive degrees of freedom across layers rather than concentrating them in a single controller [ashby1957, wiener1948cybernetics]. This also connects to cybernetic accounts of hierarchical viability and self-regulation, such as Beer’s Viable System Model [beer1995brain]. This is because weaker policies are more likely to generalise between $\mathfrak{v}$-tasks [bennett2025e].
Consider the aforementioned $\alpha_{1}$ and $\alpha_{2}$. A sufficiently weak policy at the level of individual cells is necessary for the collective to have an identity, because only the weakest policies in $\Pi_{\alpha_{1}}$ and $\Pi_{\alpha_{2}}$ will also be in $\Pi_{\omega}=\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$. A collective of cells remains cells regardless of whether I deem it to be an organ. The parts are still there. In terms of the stack, the cells $\alpha_{1}$ and $\alpha_{2}$ and the collective $\omega$ share a vocabulary. Just as the outputs of $\alpha_{1}$ can affect the inputs of $\alpha_{2}$ and vice versa, the outputs of $\omega$ determine the inputs of $\alpha_{1}$ and $\alpha_{2}$. The difference is a matter of scale. Just as w-maxing facilitates causal learning at the scale of an individual [bennett2023c], in the context of an MCA it will facilitate learning at multiple scales, because the system as a whole is just another $\mathfrak{v}$-task.
@@ 5.1 The Stack
From the enactive perspective, a human is something cells do, rather than are. An organ boils down to a collection of cells. The cells are an abstraction layer in which the collective policy of the organ is expressed, like software. However, software is just a state of hardware [bennett2024a]. Put another way, ‘organ’ is a label I ascribe to the behaviour of some cells, and ‘human’ is a label I ascribe to the behaviour of some organs. ‘x86’ is a label I ascribe to a product of human behaviour. It is a static abstraction layer an engineer has embodied in silicon and then left alone. ‘fire and forget’. Humans act as the lower layer from which x86 emerges, but once it is embodied in silicon we disconnect it from our stack. We causally ‘intervene’ in the stack. To understand the effect of this intervention, I need to abstract away the lower levels. I must “abstract away” cells and focus on the consequences of their collective behaviour.
@@
To construct layers I would start with a vocabulary $\mathfrak{v}^{i}$ at $i=0$, for example implied by a collective of cells or a lower level system. If this system exhibits goal-directed behaviour, then it is given by the $\mathfrak{v}^{i}$-task $\lambda^{i}(\mathfrak{v}^{i})$ and a policy $\pi^{i}\in\Pi_{\lambda^{i}(\mathfrak{v}^{i})}$. $\pi^{i}$ implies behaviour. I translate policy into an abstraction layer using the aforementioned $\mathfrak{f}$. $\mathfrak{f}$ is applied to $\mathrm{Ext}(\pi^{i})$ to obtain a vocabulary $\mathfrak{v}^{i+1}$, and then $\lambda^{i+1}$ is applied to $\mathfrak{v}^{i+1}$ to obtain a task $\lambda^{i+1}(\mathfrak{v^{i+1}})$. If $\Pi_{\lambda^{i+1}(\mathfrak{v^{i+1}})}\neq\emptyset$, then there exists a common purpose or “self” at this level of abstraction and I select another policy $\pi^{i+1}\in\Pi_{\lambda^{i+1}(\mathfrak{v^{i+1}})}$.
In narrative terms, adaptation at higher levels depends on adaptation at lower levels. I now state The Law of the Stack and give a brief proof sketch (the ESM provides a full proof and additional lemmas).
Formally the next theorem quantifies a delegation bottleneck, that utility at layer i+1i{+}1 is upper-bounded by the weakness (slack) of the realised policy at layer $i$.
@@
Let $\langle\lambda_{0},\lambda_{1},\dots,\lambda_{n}\rangle$ be a stack (MLA) with abstractor $\mathfrak{f}$ and state $\langle(\pi_{0},\dots,\pi_{n}),(v_{0},\dots,v_{n})\rangle$ such that $v_{i+1}=\mathfrak{f}(v_{i},\pi_{i})$ and $\pi_{i}\in\Pi_{\lambda_{i}(v_{i})}$ for all $i<n$. Fix any $i<n$, and write
@@
Assume $\Pi_{\gamma_{i+1}}\neq\emptyset$ so that $\epsilon(\gamma_{i+1})$ is well-defined. Then the achievable utility at layer i+1i\!+\!1 is upper-bounded by the weakness of the policy at layer $i$ in the following sense.
@@
Since $\Pi_{\gamma_{i+1}}\neq\emptyset$, utility $\epsilon(\gamma_{i+1})$ is attained by some correct policy at layer i+1i\!+\!1, whose extension is a subset of $L_{v_{i+1}}$, so $\epsilon(\gamma_{i+1})+|O_{\gamma_{i+1}}|\leq|L_{v_{i+1}}|$. Because $v_{i+1}$ is finite, $|L_{v_{i+1}}|\leq 2^{|v_{i+1}|}$. By construction of the abstractor, $v_{i+1}$ is the image of $\mathrm{Ext}(\pi_{i})$ under the map $o\mapsto T(o)$, hence $|v_{i+1}|\leq|\mathrm{Ext}(\pi_{i})|$. Combining yields $\epsilon(\gamma_{i+1})+|O_{\gamma_{i+1}}|\leq 2^{|\mathrm{Ext}(\pi_{i})|}$. ∎
@@ 5.2 Delegation and variational free energy
The Law of the Stack is a structural constraint. It says that degrees of freedom in higher layers is bounded by lower-layer slack. This aligns with the Free Energy Principle (FEP), which frames adaptive behaviour as variational free energy minimisation [Friston2010, friston2013, parr2024]. In the ESM (“Viability, lived history, and variational free energy”) I show that under a uniform viability prior, minimising free energy is equivalent to maximising the entropy (log-count) of viable future continuations. In Stack Theory terms, this is exactly w-maxing.
@@
The next corollary makes the delegation dependence formal. Under the uniform viability prior, abstract-layer free energy cannot be reduced below a floor set by lower-layer slack.
@@
Fix a planning horizon and let $\mu$ denote the viability task at layer i+1i{+}1. Write $O_{\mu}\subseteq L_{v_{i+1}}$ for the set of viable future continuations at that horizon. For any policy $\pi_{i+1}$ at layer i+1i{+}1 with $\Omega_{\pi_{i+1}}\neq\emptyset$, let $\Omega_{\pi_{i+1}}=\mathrm{Ext}(\pi_{i+1})\cap O_{\mu}$ be the set of viable futures compatible with $\pi_{i+1}$, and let $q_{\pi_{i+1}}$ be the uniform distribution on $\Omega_{\pi_{i+1}}$. Under the uniform viability prior assumed in the ESM, the corresponding (base-2) free energy satisfies
@@
In particular, free energy minimisation at layer i+1i{+}1 is bottlenecked by $|\mathrm{Ext}(\pi_{i})|$, so reducing $F_{2}(q_{\pi_{i+1}})$ requires sufficient lower-layer slack, i.e. a weak enough realised policy at the lower layer.
@@
Remark (accuracy and complexity). In general, variational free energy admits a decomposition into an accuracy term and a complexity term. The uniform-viability model used in Corollary 5.1 folds accuracy into hard constraints, namely the viability constraint $\omega\in\Omega$, and the history-consistency restriction $\pi\in\Pi_{\mathfrak{h}}$ within the confines of an embodied vocabulary $\mathfrak{v}$. Utility and viability constrain what counts as accurate behaviour, and delegation affects residual flexibility within that accuracy envelope.
@@ 5.3 Human Abstractions
If an abstraction layer is a window onto the world, then utility is how “good” the view is for a task. A static abstraction layer is a window that cannot move, while an MCA can change policy at lower levels, to move the window into an ideal location. When I create an ISA I insert a static abstraction layer at an intermediate level, so the vocabulary at the highest levels is limited to what I can do with those abstractions I have embodied in the ISA. In comparison, an MCA can learn policies at lower levels, to change the vocabulary of higher levels and facilitate multilayer causal learning. This shows how adaptability at higher levels demands adaptability at lower levels. I am not suggesting lower levels only optimise for utility at higher levels. After all, an increase in utility may increase the computational cost of searching for correct policies. Furthermore, a spatially extended environment constrains the size of vocabularies at all levels [bennett2024b], because bounded systems can contain only finite information [bekenstein1981]. I mean lower levels learn an inductive bias for higher levels, balancing utility with other factors in service of goals at every scale.
@@ 6 Cancer-Like Failure Modes
I have explored some advantages of bio-mimetic distribution and delegation. Are there disadvantages? Yes. A state analogous to cancer exists for distributed systems with insufficient delegation in adverse conditions. Some have argued cancer is a loss of collective identity [levin2021, davies2011]. In a biological system, when cells are isolated from the informational structure of their collective, they lose their high level identity. When that happens, they independently pursue their own lower level goals. They revert to primitive transcriptional behaviour, while the collective of which they were formerly part continues on without them. They escape from immune control mechanisms by becoming less differentiated, and closer to embryonic states.
@@
Formally, the next definition and proposition show that if boundary conditions over-constrain a distributed system so that no collective policy exists, then any realised behaviour must be splintered, because it can only satisfy a proper subset of parts.
This is a formal analogue rather than a biological claim. It captures the failure mode where the collective-policy intersection disappears while local viability persists, so continued function is only possible via coherent subcoalitions. In Stack Theory terms, this is the zero-slack limit of the delegation bottleneck quantified by Theorem 5.1. Tightening lower-layer constraints can shrink the induced higher-layer feasible set all the way to emptiness.
@@
Proof. If some $\pi$ were feasible for all parts then $\pi\in\bigcap_{j=1}^{m}\widehat{\Pi}_{\alpha_{j}}=\widehat{\Pi}_{\omega}$, contradicting $\widehat{\Pi}_{\omega}=\emptyset$. Hence every realised policy fails at least one part’s constraints, so $S(\pi)\neq\{1,\dots,m\}$ for all $\pi\in L_{v}$. If $S(\pi)\neq\emptyset$ then $\pi\in\bigcap_{j\in S(\pi)}\widehat{\Pi}_{\alpha_{j}}$, so $S(\pi)$ is a splinter. If $\widehat{\Pi}_{\alpha_{j}}\neq\emptyset$ then $\{j\}$ is a splinter since $\bigcap_{k\in\{j\}}\widehat{\Pi}_{\alpha_{k}}=\widehat{\Pi}_{\alpha_{j}}\neq\emptyset$. $\square$
Remark (integration and $\Phi_{\mathrm{IIT}}$). Proposition 6.1 characterises collective identity as the existence of a non-empty intersection of feasible policy-sets. This is a structural notion of integration. In the Stack Theoretic depiction of consciousness, a loss of collective identity would necessarily imply loss of selves (e.g. reafference) and the disintegration of phenomenal content [bennett2025thesis, bennettwelshciaunica2024, bennett2026a]. Beyond Stack Theory, Integrated Information Theory formalises integration via the quantity $\Phi_{\mathrm{IIT}}$, which decreases when a system decomposes into weakly coupled parts [tononi2004, oizumi2014]. Under that lens, splintering can be read as a loss of integration, rather than merely an over-constraint at one layer.
Remark (temporal splintering and negotiation). The set-intersection definition in Proposition 6.1 is static. Hybrid agents often face transient conflicts where $\widehat{\Pi}_{\omega}(t)=\emptyset$ for short intervals. A natural dynamic refinement is to treat each part’s feasible set as time-indexed and to model resolution as a higher-level regulatory process that reallocates variety across layers to restore a non-empty viable intersection. This is closely aligned with cybernetic treatments of adaptive organisation, including Beer’s Viable System Model and Ashby’s requisite variety principle [beer1995brain, ashby1957]. Future work may further extend this framing of cancer using the time semantics module associated with Stack Theory [bennett2026a].
@@
When $\Pi_{\lambda^{i}(v^{i})}=\emptyset$, the collective has no viable way to act as a coherent unit. Any continued function therefore requires dropping constraints, parent task splintering into ever narrower child tasks until those tasks are narrow enough to admit a policy. Part of the collective stops participating in the collective policy and reverts to a lower-level policy set with a much larger extension (more autonomy). This is a form of “selective forgetting” [bennett2023c] applied to constraints rather than data.
This is analogous to cancer. $\alpha$ is isolated and can now pursue goals which do not align with those of the collective. This suggests systems should be minimally constrained so as to ensure graceful degradation or ‘sloppy fitness’ [agosta2010], maintaining core functionality in as wide a variety of unanticipated circumstances as possible. The more difficult the environment or the more top-down the system is controlled, the more tightly constrained behaviour becomes until the system becomes over-constrained and splinters. Co-operation becomes impossible and the parts of a system start to compete. Conversely, if control of a system is delegated to the lowest levels, then adverse conditions are less likely to eliminate all correct policies. Put another way, there will be less incentive for the parts of a system to splinter, as that would not greatly increase the number of correct policies available to them. A state analogous to cancer occurs in a distributed system when there is an over-reliance on top-down causation to make the parts of a system conform to a collective identity they are ill suited to, under the circumstances.